In snapdragon plants, flower color is genetically determined as follows: 

red = rr white = ww pink = rw 


Determine the genotypic frequencies expected in a cross between 2 pink snapdragons. 
Answer #1

Pink snapdragon has the genotype “rw”

Parent                                                  Pink                       *                             Pink

                                                            rw                                                           rw

Gametes                                      r               w                                           r                   w

Fusion of these gametes will result

Pink

Pink

r

w

r

rr

rw

w

rw

ww

This cross between the two pink plant results in the formation of 3 different kind of genotypes, which are “rr”(red), “rw” (pink) and “ww” (white).

The genotypic frequency for homozygous red (rr) = 1 / 4 = 0.25 = 25%

The genotype frequency for heterozygous pink (rw) = 2 / 4 = 0.50 = 50%

The genotype frequency for homozygous white (ww) = 1 / 4 = 0.25 = 25%